If 8.42 x 10^(24)molecules of ammonium sulphate are allowed to react with excess sodium hydroxide, calculate the volume of ammonia gas produced.
1
Expert's answer
2019-05-14T03:40:01-0400

(NH4)2SO4 + 2NaOH = Na2SO4 + 2NH3 + H2O

n ((NH4)2SO4 = N/Na

n = 8,42•1024 / 6,02•1023 = 13.99 mole

n (NH3) = 27.97 mole

V (NH3) = n×Vm

V (NH3) = 27.97 × 22.4 = 626.6 litre